OpenStudy (geneticrockhopper247):

Describe the vertical asymptote and hole for the graph of x^2+x-6/x^2-9.

OpenStudy (campbell_st):

ok... you need to factor the numerator and denominator are you able to do that..?

OpenStudy (geneticrockhopper247):

I can... should I do them individually and then but them together?

OpenStudy (campbell_st):

nope just factor them and leave it in fraction form... can you show what you have... so I can check

OpenStudy (geneticrockhopper247):

(x - 2)(x + 3)/(x - 3)(x + 3)

OpenStudy (campbell_st):

ok... if you factorised you have \[\frac{(x + 3)(x -2)}{(x+3)(x-3)}\] as you can see (x+ 3) is a factor in the numerator and denominator so this will result in a hole, or point of discontinuity vertical asymptotes are the values of x that make the denominator zero... so after removing the common factor of (x + 3) your denominator is x - 3 you need to solve x = 3 = 0 to find the vertical asymptote.

OpenStudy (geneticrockhopper247):

so the vertical asymptote is x=-3?

OpenStudy (campbell_st):

so the point of discontinuity is when x + 3 = 0 ..... or x = -3 substitute this into the simplified equation of y = (x -2)/(x - 3) and you'll find the hole, or point of discontinuity occurs at (-3, 5/6) hope this helps

OpenStudy (campbell_st):

oops the you need to solve x - 3 = 0 for the vertical asymptote

OpenStudy (campbell_st):

sorry about the typo

OpenStudy (geneticrockhopper247):

can you help me with one more?

OpenStudy (campbell_st):

sorry I have to go.... but think about vertical asymptotes are the values that make the denominator zero

OpenStudy (geneticrockhopper247):

okay, thanks anyway